# JMMMU: A Japanese Massive Multi-discipline Multimodal Understanding Benchmark for Culture-aware Evaluation

### v1

@@ Abstract
Accelerating research on Large Multimodal Models (LMMs) in non-English languages is crucial for enhancing user experiences across broader populations. In this paper, we introduce JMMMU (Japanese MMMU), the first large-scale Japanese benchmark designed to evaluate LMMs on expert-level tasks based on the Japanese cultural context. To facilitate comprehensive culture-aware evaluation, JMMMU features two complementary subsets: (i) culture-agnostic (CA) subset, where the culture-independent subjects (e.g., Math) are selected and translated into Japanese, enabling one-to-one comparison with its English counterpart MMMU; and (ii) culture-specific (CS) subset, comprising newly crafted subjects that reflect Japanese cultural context. Using the CA subset, we observe performance drop in many LMMs when evaluated in Japanese, which is purely attributable to language variation. Using the CS subset, we reveal their inadequate Japanese cultural understanding. Further, by combining both subsets, we identify that some LMMs perform well on the CA subset but not on the CS subset, exposing a shallow understanding of the Japanese language that lacks depth in cultural understanding. We hope this work will not only help advance LMM performance in Japanese but also serve as a guideline to create high-standard, culturally diverse benchmarks for multilingual LMM development. The project page is https://mmmu-japanese-benchmark.github.io/JMMMU/.
@@
Shota Onoharathanks: Equal contribution.   Atsuyuki Miyai11footnotemark: 1   Yuki Imajuku11footnotemark: 1   Kazuki Egashira11footnotemark: 1   Jeonghun Baek11footnotemark: 1 Xiang Yue    Graham Neubig   Kiyoharu Aizawa The University of Tokyo    Carnegie Mellon University
@@ 1 Introduction
In recent years, large language models (LLMs) have revolutionized the field of language processing (Chen et al., 2023a; vic, 2023; Touvron et al., 2023; Wei et al., 2023). Building on the success of LLMs, large multimodal models (LMMs) have demonstrated remarkable performance across tasks ranging from common sense reasoning to domain-specific, expert-level challenges (Antol et al., 2015; Liu et al., 2023a, 2024c; Yue et al., 2024). As their capabilities grow, the need for robust criteria to evaluate LMMs has become increasingly important, highlighting the role of comprehensive benchmarks in assessing the full scope of their abilities.
However, current benchmarks focus primarily on performance in English Liu et al. (2024c); Yue et al. (2024); Li et al. (2024b); Liu et al. (2023b); Yu et al. (2024); Fu et al. (2024), with less emphasis on evaluation in other languages. Given that LMMs are widely used across diverse languages, it is imperative to evaluate their performance beyond English. Additionally, such multilingual evaluations should actively involve contributions from diverse communities, ensuring that the associated cultural contexts are appropriately considered.
In this paper, we introduce JMMMU (Japanese MMMU), the first benchmark designed to evaluate LMMs on extensive, multi-disciplinary tasks in Japanese that require college-level subject knowledge, deliberate reasoning, and cultural understanding. The overview of JMMMU is shown in Figure 1. JMMMU draws inspiration from the well-established MMMU (Yue et al., 2024) and expands existing culture-aware Japanese benchmarks Inoue et al. (2024b); SakanaAI (2024c) by over 10 times, with 1,320 questions using 1,118 images, covering a diverse range of subjects.
@@
Evaluating 15 open-source LMMs and three advanced proprietary LMMs on JMMMU, our key findings are summarized as follows:
@@ 2 Related Work
Following the success of large language models (LLMs), many LMMs have been developed with improved knowledge and instruction-following capabilities (Liu et al., 2023b, 2024a, 2024b; Li et al., 2024a; Ye et al., 2024; Zhao et al., 2023; Li et al., 2023; Monajatipoor et al., 2023; Zhao et al., 2024). However, the progress of these models is typically evaluated on English benchmarks (Yue et al., 2024; Liu et al., 2024c). Therefore, a significant challenge remains in accurately evaluating the capabilities of other languages, highlighting the need for non-English benchmarks.
@@
Among various recent benchmarks (Li et al., 2024b; Liu et al., 2023b, 2024c; Lu et al., 2024; Yue et al., 2024; Miyai et al., 2024), MMMU (Yue et al., 2024) is the most widely used to measure the advancements of cutting-edge LMMs. MMMU requires advanced university-level knowledge and reasoning across a broader range of subjects, enabling a more comprehensive and expert-level evaluation. Subsequently, CMMMU (Zhang et al., 2024) has been proposed as its Chinese counterpart. While CMMMU comprises entirely new culture-specific questions, our JMMMU has not only culture-specific subjects but also translation-based culture-agnostic subjects, facilitating one-to-one comparisons between English and Japanese using the exact same questions. In line with multilingual ability evaluation, several VQA benchmarks have been proposed (Gao et al., 2015; Changpinyo et al., 2022; Gupta et al., 2020; Liu et al., 2021; Pfeiffer et al., 2021; Tang et al., 2024; Romero et al., 2024). However, unlike the MMMU series, their primary focus is on daily knowledge, (e.g., Pop Culture, Sports in CVQA (Romero et al., 2024)), still leaving the multilingual expert-level reasoning skills as an important direction for future work.
@@ 3.2 Data Curation Process
The images are primarily sourced from Wikimedia Commons [^1] https://commons.wikimedia.org/, ensuring that all selected images are available under licenses suitable for public release. In crafting questions, we aimed to keep the text as simple as possible and ensure that no options stand out, making it hard to guess the correct choice without referring to the image.
@@ 4.1 Setup
  • Open-source LMMs: LLaVA-OneVision 0.5B & 7B (Li et al., 2024a), LLaVA1.6-13B & 34B (Liu et al., 2024b), Phi-3 & 3.5 Vision (Abdin et al., 2024), InternVL2-2B & 8B (Chen et al., 2023b), xGen-MM (Xue et al., 2024), Idefics2-8B (Laurençon et al., 2024b), Idefics3-8B (Laurençon et al., 2024a), CogVLM2-19B (Hong et al., 2024), and Mantis-8B (Jiang et al., 2024).
In Appendix A, we provide further details of these models, with a particular focus on Japanese language support.
@@
The evaluation method is based on the setup in MMMU (Yue et al., 2024). Prompts are translated as follows: for multiple-choice questions, 与えられた選択肢の中から最も適切な回答のアルファベットを直接記入してください。 (Answer with the option’s letter from the given choices directly.) ; and for open-ended questions, 質問に対する回答を単語や短いフレーズで記入してください。 (Answer the question using a single word or phrase.).
@@ 4.2 Main Result
Table 2 demonstrates the evaluation results on our JMMMU benchmark. We provide the average scores across all subjects, culture-agnostic (CA) subjects, and culture-specific (CS) subjects, as well as scores on individual subjects. For comparison, we also provide the performance on CA suset in English CA (EN) . Note that CA (EN) is often smaller than the overall average of MMMU given by Yue et al. (2024) because subjects selected as CA are relatively difficult among all subjects in MMMU as it often requires stronger reasoning capabilities (e.g., Math).
@@
First, as a general trend, the score on the CA subset is significantly lower than its English counterpart (CA (EN) in Table 2) with an average drop of 4.5%. This indicates that, even for the same questions, many models perform worse when asked in Japanese. Second, despite such a general trend, Japanese-made LMMs (i.e., LLaVA CALM2 and EvoVLM JP v2) face a minimal drop (up to 1.7 %), which implies that incorporating the Japanese dataset successfully mitigates the performance gap between English and Japanese.
Figure 3: Score correlation between subsets. While proprietary models ($\blacksquare$) perform the best on both subsets, Japanese LMMs ($\bigstar$) perform remarkably high on CS subset compared to models that perform similarly on CA subset.
@@
Figure 3 demonstrates the correlation between the scores on the CA and CS subjects. The Japanese LMMs, LLaVA CALM2 and EvoVLM JP v2, show higher scores on CS subjects compared to other models that perform similarly on CA subjects. This strongly indicates their proficiency in CS subjects. On the other hand, however, compared to stronger models such as InternVL2-8b, LLaVA1.6-34b, and LLaVA-OV-7b, the Japanese LMMs show lower scores on CA subjects, suggesting room for improvement in their general reasoning and problem-solving capabilities in culture-agnostic context.
@@ 5.2 Errors in Culture-agnostic Subjects
JMMMU shares 720 culture-agnostic questions with MMMU, which allows us to compare the output one by one. Using these questions, we evaluate how translation affects model performance. Taking GPT-4o as an example, we classify the responses into four categories based on whether they are correct or incorrect in each language. LABEL:fig:CA_error presents the results before and after translation. The results on the other models are provided in Section B.1
@@
While GPT-4o performs similarly in both languages on culture-agnostic split (only 0.3% difference in Table 2), we have found that there are a significant amount (28.6 %) of questions to which it answered correctly only in either one of the languages. We now investigate this phenomenon. For questions answered correctly only in English (orange in LABEL:subfig:ca_error), we observe simple performance degradation after translation. In contrast, we have found some distinctive examples in the opposite case (yellow). In an example of LABEL:subfig:before_and_after_translation, GPT-4o outputs only the direct answer in English, whereas in Japanese, the model includes the reasoning process in its response although the model is instructed to generate the choice directly by using the prompt in Section 4.1. For a fair comparison with MMMU (Yue et al., 2024), we count a response to be correct as far as the model’s response is accurate and can be parsed by a rule-based algorithm, regardless of its instruction-following ability. As a result, the scores can sometimes be counterintuitively overestimated due to the lack of instruction following skills in Japanese. While the primary focus of JMMMU is on evaluating expert knowledge and supporting the improvement of such capabilities, our findings highlight a crucial direction for future work: measuring and enhancing instruction-following ability in non-English languages.
@@ C.2 Evaluation Protocol
While the models are instructed to answer their choice directly, they often generate some contextual information or unnecessary symbols. To tackle this point, following MMMU (Yue et al., 2024), we extract an answer from the model response with a rule-based method. For multiple-choice questions, this parser can extract the model’s choice even when the choice is surrounded by some symbol (e.g., ’(A)’, ’A.’, ’A ’) or by text.
@@ Appendix D Annotation Instruction
The document containing the instructions presented to the annotators is shown in Figure B.
Figure B: Annotation Instruction. Annotators were provided with the Japanese version of this instruction.
@@ Appendix E Examples
We provide sample questions from culture-agnostic subset in Figure C, and questions from culture-specific subset in Figure D
Figure C: Examples in culture-agnostic subjects. Some images that contain English are translated.
Figure D: Examples in culture-specific subjects. The questions are created by Japanese native speakers and requires knowledge of Japanese culture.

### v2

@@ Abstract
Accelerating research on Large Multimodal Models (LMMs) in non-English languages is crucial for enhancing user experiences across broader populations. In this paper, we introduce JMMMU (Japanese MMMU), the first large-scale Japanese benchmark designed to evaluate LMMs on expert-level tasks based on the Japanese cultural context. To facilitate comprehensive culture-aware evaluation, JMMMU features two complementary subsets: (i) culture-agnostic (CA) subset, where the culture-independent subjects (e.g., Math) are selected and translated into Japanese, enabling one-to-one comparison with its English counterpart MMMU; and (ii) culture-specific (CS) subset, comprising newly crafted subjects that reflect Japanese cultural context. Using the CA subset, we observe performance drop in many LMMs when evaluated in Japanese, which is purely attributable to language variation. Using the CS subset, we reveal their inadequate Japanese cultural understanding. Further, by combining both subsets, we identify that some LMMs perform well on the CA subset but not on the CS subset, exposing a shallow understanding of the Japanese language that lacks depth in cultural understanding. We hope this work will not only help advance LMM performance in Japanese but also serve as a guideline to create high-standard, culturally diverse benchmarks for multilingual LMM development.
@@
Shota Onohara1thanks: Equal contribution. Atsuyuki Miyai [^1] footnotemark: 1 Yuki Imajuku [^1] footnotemark: 1 Kazuki Egashira [^1] footnotemark: 1 Jeonghun Baek [^1] footnotemark: 1 Xiang Yue2     Graham Neubig2     Kiyoharu Aizawa1 1The University of Tokyo     2Carnegie Mellon University
@@ 1 Introduction
In recent years, large language models (LLMs) have revolutionized the field of language processing (Chen et al., 2023a; vic, 2023; Touvron et al., 2023; Wei et al., 2023). Building on the success of LLMs, large multimodal models (LMMs) have demonstrated remarkable performance across tasks ranging from common sense reasoning to domain-specific, expert-level challenges (Antol et al., 2015; Liu et al., 2023a, 2024c; Yue et al., 2024a). As their capabilities grow, the need for robust criteria to evaluate LMMs has become increasingly important, highlighting the role of comprehensive benchmarks in assessing the full scope of their abilities.
However, current benchmarks focus primarily on performance in English Liu et al. (2024c); Yue et al. (2024a); Li et al. (2024b); Liu et al. (2023b); Yu et al. (2024); Fu et al. (2024), with less emphasis on evaluation in other languages. Given that LMMs are widely used across diverse languages, it is imperative to evaluate their performance beyond English. Additionally, such multilingual evaluations should actively involve contributions from diverse communities, ensuring that the associated cultural contexts are appropriately considered.
In this paper, we introduce JMMMU (Japanese MMMU), the first benchmark designed to evaluate LMMs on extensive, multi-disciplinary tasks in Japanese that require college-level subject knowledge, deliberate reasoning, and cultural understanding. The overview of JMMMU is shown in Figure 1. JMMMU draws inspiration from the well-established MMMU (Yue et al., 2024a) and expands existing culture-aware Japanese benchmarks Inoue et al. (2024b); SakanaAI (2024c) by over 10 times, with 1,320 questions using 1,118 images, covering a diverse range of subjects.
@@
Evaluating 16 open-source LMMs and three advanced proprietary LMMs on JMMMU, our key findings are summarized as follows:
@@ 2 Related Work
Following the success of large language models (LLMs), many LMMs have been developed with improved knowledge and instruction-following capabilities (Liu et al., 2023b, 2024a, 2024b; Li et al., 2024a; Ye et al., 2024; Zhao et al., 2023; Li et al., 2023; Monajatipoor et al., 2023; Zhao et al., 2024). However, the progress of these models is typically evaluated on English benchmarks (Yue et al., 2024a; Liu et al., 2024c). For example, there is a study measuring LMM performance in Japanese—linguistically and culturally distinct from English—by generating recipe texts from food photos. Interestingly, these experiments found that existing proprietary LMMs sometimes perform worse than smaller open-source models fine-tuned on Japanese recipe data, suggesting that these proprietary models may lack sufficient knowledge of Japanese foods(Imajuku et al., 2025). Therefore, a significant challenge remains in accurately evaluating the capabilities of other languages, highlighting the need for non-English benchmarks.
@@
Among various recent benchmarks (Li et al., 2024b; Liu et al., 2023b, 2024c; Lu et al., 2024; Yue et al., 2024a; Miyai et al., 2024), MMMU (Yue et al., 2024a) is the most widely used to measure the advancements of cutting-edge LMMs. MMMU requires advanced university-level knowledge and reasoning across a broader range of subjects, enabling a more comprehensive and expert-level evaluation. Subsequently, CMMMU (Zhang et al., 2024a) has been proposed as its Chinese counterpart. While CMMMU comprises entirely new culture-specific questions, our JMMMU has not only culture-specific subjects but also translation-based culture-agnostic subjects, facilitating one-to-one comparisons between English and Japanese using the exact same questions. In line with multilingual ability evaluation, several VQA benchmarks have been proposed (Gao et al., 2015; Changpinyo et al., 2022; Gupta et al., 2020; Liu et al., 2021; Pfeiffer et al., 2021; Tang et al., 2024; Romero et al., 2024). However, unlike the MMMU series, their primary focus is on daily knowledge, (e.g., Pop Culture, Sports in CVQA (Romero et al., 2024)), still leaving the multilingual expert-level reasoning skills as an important direction for future work.
@@ 3.2 Data Curation Process
We carefully designed the subject selection process to include only standard fields of study at the Japanese university level. The images are primarily sourced from Wikimedia Commons [^1] https://commons.wikimedia.org/, ensuring that all selected images are available under licenses suitable for public release. In crafting questions, we aimed to keep the text as simple as possible and ensure that no options stand out, making it hard to guess the correct choice without referring to the image.
@@ 4.1 Setup
  • Open-source LMMs: LLaVA-OneVision 0.5B & 7B (Li et al., 2024a), LLaVA1.6-13B & 34B (Liu et al., 2024b), Phi-3 & 3.5 Vision (Abdin et al., 2024), InternVL2-2B & 8B (Chen et al., 2023b), xGen-MM (Xue et al., 2024), Idefics2-8B (Laurençon et al., 2024b), Idefics3-8B (Laurençon et al., 2024a), CogVLM2-19B (Hong et al., 2024), Mantis-8B (Jiang et al., 2024), and Pangea-7B (Yue et al., 2024b).
We run all experiments with LMMs-Eval Zhang et al. (2024b). In Appendix A, we provide further details of these models, with a particular focus on Japanese language support.
@@
The evaluation method is based on the setup in MMMU (Yue et al., 2024a). Prompts are translated as follows: for multiple-choice questions, 与えられた選択肢の中から最も適切な回答のアルファベットを直接記入してください。 (Answer with the option’s letter from the given choices directly.) ; and for open-ended questions, 質問に対する回答を単語や短いフレーズで記入してください。 (Answer the question using a single word or phrase.).
@@ 4.2 Main Result
Table 2 demonstrates the evaluation results on our JMMMU benchmark. We provide the average scores across all subjects, culture-agnostic (CA) subjects, and culture-specific (CS) subjects, as well as scores on individual subjects. For comparison, we also provide the performance on CA suset in English CA (EN) . Note that CA (EN) is often smaller than the overall average of MMMU given by Yue et al. (2024a) because subjects selected as CA are relatively difficult among all subjects in MMMU as it often requires stronger reasoning capabilities (e.g., Math).
@@
First, as a general trend, the score on the CA subset is significantly lower than its English counterpart (CA (EN) in Table 2) with an average drop of 4.3%. This indicates that, even for the same questions, many models perform worse when asked in Japanese. Second, despite such a general trend, Japanese-made LMMs (i.e., LLaVA CALM2 and EvoVLM JP v2) face a minimal drop (up to 1.7 %), which implies that incorporating the Japanese dataset successfully mitigates the performance gap between English and Japanese.
Figure 3: Score correlation between subsets. While proprietary models ($\blacksquare$) perform the best on both subsets, Japanese LMMs ($\bigstar$) and Pangea ($\blacklozenge$, a culture-aware multilingual LMM) perform remarkably high on CS subset compared to models that perform similarly on CA subset.
@@
Figure 3 demonstrates the correlation between the scores on the CA and CS subjects. The Japanese LMMs (LLaVA CALM2 and EvoVLM JP v2) and Pangea (a culture-aware multilingual LMM) show higher scores on CS subjects compared to other models that perform similarly on CA subjects. This strongly indicates their proficiency in CS subjects. As shown in appendix A, Japanese LMMs uses not only translated corpora but also texts originally written in Japanese. This may contribute to their performance in CS subjects. On the other hand, however, compared to stronger models such as InternVL2-8b, LLaVA1.6-34b, and LLaVA-OV-7b, the Japanese LMMs show lower scores on CA subjects, suggesting room for improvement in their general reasoning and problem-solving capabilities in culture-agnostic context.
@@ 5.2 Errors in Culture-agnostic Subjects
JMMMU shares 600 culture-agnostic questions with MMMU, which allows us to compare the output one by one. Using these questions, we evaluate how translation affects model performance. Taking GPT-4o as an example, we classify the responses into four categories based on whether they are correct or incorrect in each language. LABEL:fig:CA_error presents the results before and after translation. The results on the other models are provided in Section B.1
@@
While GPT-4o performs similarly in both languages on culture-agnostic split (only 0.3% difference in Table 2), we have found that there are a significant amount (28.6 %) of questions to which it answered correctly only in either one of the languages. We now investigate this phenomenon. For questions answered correctly only in English (orange in LABEL:subfig:ca_error), we observe simple performance degradation after translation. In contrast, we have found some distinctive examples in the opposite case (yellow). In an example of LABEL:subfig:before_and_after_translation, GPT-4o outputs only the direct answer in English, whereas in Japanese, the model includes the reasoning process in its response although the model is instructed to generate the choice directly by using the prompt in Section 4.1. Approximately 70% of the cases marked in yellow exhibit this tendency to include the reasoning process only in Japanese. For a fair comparison with MMMU (Yue et al., 2024a), we count a response to be correct as far as the model’s response is accurate and can be parsed by a rule-based algorithm, regardless of its instruction-following ability. As a result, the scores can sometimes be counterintuitively overestimated due to the lack of instruction following skills in Japanese. While the primary focus of JMMMU is on evaluating expert knowledge and supporting the improvement of such capabilities, our findings highlight a crucial direction for future work: measuring and enhancing instruction-following ability in non-English languages.
@@ Appendix A LMMs’ Japanese Support
Pangea is designed as a multilingual LMM. This model is trained on 6 million instruction datasets across 39 languages, including Japanese.
@@ B.3 Score Correlation between languages
B.4 Chain of Thought Performance
We investigate the effect of Chain of Thought (CoT) prompting on performance enhancement. The prompts are derived from the CoT prompts utilized in MMMU-Pro Yue et al. (2024a), translated into Japanese for this study. The prompts are translated as follows: for multiple-choice questions, 「与えられた選択肢の中から最も適切な回答のアルファベットを直接記入してください。1歩ずつ考えてください。そして最後に選択肢の答えを次の形式で答えてください:答え: \$LETTER(かぎかっこなし)ここで LETTER は選択肢のいずれかです。」 (“Answer with the option’s letter from the given choices directly. The last line of your response should be of the following format: ‘Answer: \$LETTER’ (without quotes) where LETTER is one of the options. Think step by step before answering.”); and for open-ended questions, 「質問に対する回答を単語や短いフレーズで記入してください。1歩ずつ考えてください。そして最後の行に答えのみを出力してください。」 (“Answer the question using a single word or phrase. Think step by step. Finally, output only the answer on the last line.”) Figure B illustrates the impact of CoT prompting.
Figure B: Impact of CoT prompting in JMMMU. Only GPT-4o demonstrates score improvements.
Figure C: Impact of English direct prompting in JMMMU.
Although GPT-4o achieves a significant improvement in performance, the open-source models show only marginal gains, with a maximum increase of $+1.5\%$. One possible explanation, as also mentioned in MMMU-Pro, is the limited instruction-following ability of these models. In addition to this, their difficulty in generating long outputs in Japanese further hinders their performance. This limitation impacts the models’ ability to produce detailed reasoning processes, leading to instances where the correct answer cannot be reached.
B.5 English Direct Prompt
In LLMs, it has been shown that performance decreases when using prompts in languages other than English compared to English (Shi et al., 2023).
To investigate reasoning abilities across different languages in JMMMU, we evaluate the models with English direct prompts. English Direct Prompts refer to cases where the question remains in Japanese, but the instruction is given in English. Direct Prompts are defined as follows: for multiple-choice questions, “Answer with the option’s letter from the given choices directly.”; and for open-ended questions, “Answer the question using a single word or phrase.”
We show the results in fig. C. Notable improvements in scores are observed in Pangea’s CA subjects and GPT-4o’s CS subjects.
A detailed analysis of the model outputs reveals that Pangea exhibits significant deficiencies in instruction-following ability when prompted in Japanese, frequently failing to provide answers in correct formats. In contrast, the underlying cause of the substantial score increase in GPT-4o’s CS subjects remains unclear and requires further investigation.
@@ C.2 Evaluation Protocol
While the models are instructed to answer their choice directly, they often generate some contextual information or unnecessary symbols. To tackle this point, following MMMU (Yue et al., 2024a), we extract an answer from the model response with a rule-based method. For multiple-choice questions, this parser can extract the model’s choice even when the choice is surrounded by some symbol (e.g., ’(A)’, ’A.’, ’A ’) or by text.
@@ Appendix D Annotation Instruction
The document containing the instructions presented to the annotators is shown in Figure D.
Figure D: Annotation Instruction. Annotators were provided with the Japanese version of this instruction.
@@ Appendix E Examples
We provide sample questions from culture-agnostic subset in Figure E, and questions from culture-specific subset in Figure F
Figure E: Examples in culture-agnostic subjects. Some images that contain English are translated.
Figure F: Examples in culture-specific subjects. The questions are created by Japanese native speakers and requires knowledge of Japanese culture.